OpenStudy (anonymous):

Hallie is trying to win the grand prize on a game show. Should she try her luck by spinning a wheel with 6 equal sections labeled 1 to 6 and hope she gets a 5, or should she roll two number cubes and hope she gets the same number on both cubes? Explain.

OpenStudy (jim766):

what is the probability of winning at the wheel?

OpenStudy (anonymous):

Either method is equally likely. They both will produce a success at: Probability = 1/6 The spin will give success at 1 out of 6 The dice roll will give success at 6 out of 36 -> 1 out of 6 They are equal as far as probability goes.

OpenStudy (jim766):

the probability of rolling a 6 on a die is 1/6 the probability of rolling a 6 on a 2nd die is 1/6 isnt probability of rolling both 1/6 * 1/6

OpenStudy (anonymous):

@Jim766 Yes it is

OpenStudy (anonymous):

1/6*1/6 = 1/36

OpenStudy (anonymous):

A little bit more about the dice rolls: There are 36 different dice rolls: any of 6 #'s on 1 die multiplied by any of 6 #'s on the other die. 6 x 6 = 36. And there are only 6 different rolls where both #'s are the same: 1,1 2,2 3,3 4,4 5,5 6,6

OpenStudy (jim766):

so which is better probability 1/6 or 1/36?

OpenStudy (anonymous):

Spinnig wheel P(of getting a 5) 1/6 Number cube = 6 sides 2 number cubes = 6*6=36 and probaibily is 6 of getting same number 6/36 u can divide by 6 to get 1/6 same chance

OpenStudy (anonymous):

That's why the P(2 same #'s) = 1/6 -> 6 out of 36 = 1/6 And as for getting a "5" when spinning: 1/6 That's why the two activities give the same probability.

OpenStudy (anonymous):

@Jim766 ,the dice roll situation is not calling for both #'s to be a "5". It is calling for both numbers to be the same, so it's not 1/36, it's 6/36 = 1/6

OpenStudy (jim766):

ahh so the difference is that they were not specific about what number had to be rolled...

OpenStudy (anonymous):

Actually they were specific in not requiring any particular number. That lack of that particular requirement is in itself a "specificity".

OpenStudy (anonymous):

Therefore, she could pick either one and get the same chance of winning no matter which one she picks?

OpenStudy (anonymous):

That's right!

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

uw! Good job everyone!

OpenStudy (anonymous):

All good now, @mangorox ?

OpenStudy (anonymous):

Yes, thanks @tcarroll010 !

OpenStudy (anonymous):

Good luck in all of your studies! @mangorox